One monkey climbs a poll at the rate of 6mts/min and fell down 3mts in the alternately. Length of the poll is 60 mts , how much time it will take to reach the top? a. 31 b.33 c.37 d.40

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Effective distance covered in 1 min=6-3=3m

Hence total time taken will be 30+3=33min

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi